The average hourly wage of employees of a certain company is $9.83. Assume the variable is normally distributed. If the standard deviation is $4.58, find the probability that a randomly selected employee earns less than $5.43.
0.313 = 31.3%
0.168 = 16.8%
0.436 = 43.6%
0.345 = 34.5%

: The average hourly wage of employees of a certain company is $9.83. Assume the variable is normally distributed. If the standard deviation is $4.58, find the probability that a randomly selected employee earns less than $5.43.
0.313 = 31.3%
0.168 = 16.8%
0.436 = 43.6%
0.345 = 34.5%


Answer by stanbon(18984)

The average hourly wage of employees of a certain company is $9.83. Assume the variable is normally distributed. If the standard deviation is $4.58, find the probability that a randomly selected employee earns less than $5.43.
-------------------
Find the z-score of 5.43
z(5.43) = (5.43-9.83)/4.58 = -0.960699
-------
Then P(x < 5.43) = P(z < -0.960699) = 0.1684 = 16.84%
======================================
Cheers,
Stan H.
=================
0.313 = 31.3%
0.168 = 16.8%
0.436 = 43.6%
0.345 = 34.5%